### v1

# Destruction of ’Peas in a Pod?’ A Candidate Multi-planet System Around the Nearby, Bright Star, HD208487

Abstract
Context. The search for extrasolar planets (aka exoplanets) is continuously intensifying as new and improved instrumentation provides more precise data. Discovering new exoplanets orbiting bright stars in the local solar neighborhood is particularly important, as they tend to provide the best targets for future follow-up studies.
Aims. We re-investigate the HD208487 star system to test the reality of the proposed HD208487c world. We also search for additional companions using applied Bayesian statistics and 15+ years of new radial velocity (RV) data from the HARPS and the PFS instruments that were taken post-discovery of HD208487b, as part of our continued study of bright Sun-like stars within 50 pc of the Sun.
Methods. The RV data was analyzed with Generalized Lomb-Scargle Periodograms, followed by Bayesian analysis techniques using the EMPEROR code. We also scrutinised various stellar activity indices to search for any corresponding peaks in the power spectra, correlations with the RV measurements, or significant signals from a Bayesian analysis methodology. Finally, photometric data was also checked to test for any transits or possible activity manifestations that could lead to possible false RV signals or excess noise.
[…]
Conclusions. We have uncovered a candidate three planet system that would consist of an inner gas giant, a central Saturn and an outer super-Neptune/sub-Saturn. Extensive analysis of both photometric and spectroscopic data as activity proxies strongly supports the planetary system hypothesis, yet more long-term RV data would help add more statistical weight to the reality of candidate planets c and d.Assuming our model best represents reality, a dynamical analysis suggests that gravitational scattering of an initially ordered, equally-spaced system in a long resonant chain of six Neptunes can explain the current architecture of HD208487- a moderately-eccentric, inner massive planet with an outer nearly resonant ($P_{d}/P_{c}=1.495$) gas giant and super-Neptune. More RVs may also shed light on the reality of a fourth Doppler signal uncovered in the data that sits close to the 2:1 period-ratio with signal of HD208487c.
Key Words.:
Exoplanets – HD208487 – Neptune – Multiplanetary Systems – Radial Velocity Method
1 Introduction
In 1995, Mayor and Queloz discovered 51 Peg b, the first confirmed exoplanet around a main sequence star, using the Radial Velocity (RV) method (Mayor & Queloz, 1995). This planetary detection represented a paradigm shift and launched the exoplanet field on the path towards its current sophisticated state. Specifically, their discovery proved giant planets could be detected with the technology at the time, resulting in a huge push to find more new worlds, such as 70 Vir b (Marcy & Butler, 1996), and 47 UMa b (Butler & Marcy, 1996). Subsequently, more advanced telescopes with much higher resolution and precision were developed, allowing the detection of less massive planets, including, for example, the super-Earths GJ667C b, c, and d (Anglada-Escudé et al., 2012, 2013). Additionally, astronomers searched for planets in the Habitable Zone (HZ), where liquid water could exist on a planet’s surface. Santos et al. (2001) discovered HD28185, which Jones et al. (2006) confirmed to be the first planet entirely within the HZ. Eventually, Quintana et al. (2014) discovered Kepler 186f, the first (of many) ‘Earth-like’ planets orbiting its star in the habitable zone. As more and more discoveries accumulated, multi-planetary solar systems were discovered, including Upsilon Andromodae b, c, and d (Lissauer, 1999). Further, there have now been many detections of multi-planetary systems in which two or more planets are locked in a mean-motion orbital resonance (MMR) (Popova & Shevchenko, 2014; Beaugé et al., 2006; Ramos et al., 2017), such as the discovery of TOI-216b and c, two warm, large exoplanets in a 2:1 MMR (Dawson et al., 2019).
As the body of planetary systems with confirmed or candidate planets increases, it remains extremely important to revisit these systems to ensure the most accurate planetary parameters are recorded. It is not uncommon for a planet detection to later be contested due to future analyses with more data or differing methods producing inconsistent results (Jones & Jenkins, 2014; Wittenmyer et al., 2013, 2019a, 2019b; Jenkins et al., 2013; Jenkins & Tuomi, 2014; Santos et al., 2014). For example, it is well-known that stellar activity can create periodic RV signals (Oshagh et al., 2017; Costes et al., 2021). Magnetic activity, photometric variations, stellar rotation cycles, or long-term magnetic cycles can imitate planetary RV signals (Dumusque et al., 2011; Lovis et al., 2011). Therefore, examining available stellar indices, photometric data, and any correlations with the RV data remains an important step in verifying true planetary RV signals (e.g. see D´ıaz et al., 2018). On the other hand, re-examinations of previously studied star systems frequently lead to new planetary detections (Mayor et al., 2009; Vogt et al., 2010; Tuomi et al., 2013; Orell-Miquel et al., 2023).
1.1 Planet System Configurations
The scientific community places a special emphasis on the search for exoplanets in nearby solar systems, a case in point being the discovery of Proxima b, a terrestrial world orbiting within the HZ of our nearest neighbour star, Proxima Centauri (Anglada-Escudé et al., 2016), since these offer immense hope to advance our understanding of planet formation and evolution. HD208487 serves as a particularly interesting system to re-examine, since it is one of the nearest ($\sim 45$pc) and brightest ($V=7.47$) stars in the southern sky, and it already is known to be orbited by one planet. Additionally, a second, longer period signal was detected (Gregory, 2005, 2007), though later disregarded as likely activity-induced (Goździewski & Migaszewski, 2006; Wright et al., 2007). We now have 184 new RV observations spanning more than 15 years since Tinney et al. (2005) published this original discovery of HD208487b.
Of the nearly 900 multiplanetary systems [^1] The Extrasolar Planets Encyclopaedia (http://exoplanet.eu), only a handful of these exhibit similarities to our own Solar System. Weiss et al. (2018) analysed the architectures of the Kepler sample of multi-planet systems, unveiling a phenomenon denoted as ‘peas in a pod.’ Within a multi-planet system, each planet’s size is strongly correlated with that of its neighboring planets. In systems boasting three or more transiting planets, an additional pattern emerges: a systematic regularity in the spacing of planets. Frequently, these systems are densely populated with planets near their stability thresholds. Extensive research on the dynamic evolution and stability indicates that gravitational scattering plays a crucial role in shaping the final architecture of such multi-planetary systems (e.g., Chatterjee et al., 2008; Johansen et al., 2012; Izidoro et al., 2017; Wu & He, 2023; Ghosh & Chatterjee, 2024). This process triggers perturbations in planetary orbits, giving rise to close encounters between planets. Ultimately, this can lead to either collisions (whether between two planets or between a planet and its star) or the ejection of planets from the system. The former is more prevalent at smaller semi-major axes (see, e.g., Petrovich et al., 2014). This observed correlation between planet size and spacing underscores the pivotal role of dynamics in shaping the final sizes and orbital separations of planets within these systems.
In this paper, we will discuss how our new analysis has lead to the statistical confirmation and discovery of two new planets in this system, HD208487c and HD208487d. The rest of the paper will follow the ensuing structure: §2 will discuss the RV observations in detail, including a breakdown of the measurements by telescope. In §3 we present updated stellar parameters for HD208487 using the SPECIES and ARIADNE codes. We begin the analysis of the RV dataset by observing the GLS periodograms in §4, while §5 details the detection of the three planets with the more rigorous Bayesian analysis. Next, we highlight the search for correlations between stellar activity indices and the RVs in §6. §7 describes the analysis of ASAS photometric data and the fitting of stellar cycles. Finally, we discuss the results and their implications in §8, followed by the conclusions in §9.
2 RV Observations
We used a total of 219 high-precision Doppler spectroscopic RV measurements of HD208487 spanning from JD 2451034.18, August 8, 1998, to JD 2459508.64, October 21, 2021 for a total baseline of 8474 days (23.2 years), as seen in Figure 1. The data comes from the University College London Échelle Spectrograph (UCLES) instrument installed on the 3.9m Anglo-Australian Telescope at the Siding Spring Observatory (SSO), the High Accuracy Radial Velocity Planet Searcher (HARPS) installed on the 3.6m ESO Telescope (before and after the fibre upgrade in 2015) at La Silla Observatory, and the Carnegie Planet Finder Spectograph (PFS) mounted on the 6.5m Magellan II Telescope at Las Campanas Observatory.
Figure 1: All 219 mean-subtracted radial velocity measurements of HD208487, including 48 UCLES (grey), 20 HARPS-TERRA (red), 106 HARPS-TERRA2 (olive), and 45 PFS (teal) observations.
2.1 3.9m University College London Échelle Spectrograph (UCLES)
A total of 48 observations of HD208487 were collected using UCLES at the 3.9m Anglo-Australian Telescope (AAT) between JD 2451034.18 (August 8, 1998) to JD 2456969.96 (November 8, 2014) as part of the Anglo-Australian Planet Search (AAPS). UCLES is a cross-dispersed echelle spectrograph, which uses an $\text{I}_{2}$ absorption cell to derive RVs at the $\sim 3$m/s precision level (Butler et al., 2001). In addition to bias subtraction and flatfielding, the $\text{I}_{2}$ method utilizes a point spread function (PSF) to properly wavelength calibrate the spectra. The known $\text{I}_{2}$ spectral features are used to calibrate the wavelengths of the stellar spectra and to model the PSF. Next, the stellar spectrum is deconvolved from the densely-packed $\text{I}_{2}$ features to successfully extract the correct Doppler shifts found in the observed stellar spectra, which enables the calculation of long-term stable, precise RVs (Butler et al., 1996). UCLES offers a resolution of 40,000-120,000 depending on the slit width and CCD binning (Diego et al., 1990; Tinney et al., 2005). The precision of UCLES data is between $2-3\text{ms}^{-1}$ (Butler et al., 2006). These data were obtained from the NASA Exoplanet Archive [^2] https://exoplanetarchive.ipac.caltech.edu/ from (Butler et al., 2006) and from newly measured RV data calculated for this work (Butler private communication). These data will henceforth be referred to as UCLES.
2.2 3.6m High Accuracy Radial velocity Planet Searcher (HARPS)
HARPS is a stabilized, high-precision fiber-fed echelle spectrograph, contained in a vacuum cell (variations maintained $\leq 0.01$mbar) surrounded by a thermally stablized enclosure in order to avoid temperature fluctuations ($\leq 0.01$K) affecting the RV measurements (see Lovis et al., 2006). This instrument utilizes simultaneous observations through two fibers, one typically used to image the target and the other a Th-Ar reference spectrum to maintain high wavelength stability and monitoring over the long-term. With a fibre aperture on the sky of 1”, HARPS maintains a resolving power of around 115,000. HARPS achieves an overall radial velocity accuracy of $\sim 1$m/s (Mayor et al., 2003; Figueira et al., 2010).
2.2.1 TERRA
Using publicly available data from HARPS obtained from the ESO HARPS archive (PIs: Mayor, M; Kuerster, M; Trifonov, T), we found 20 observations between JD 2453265.69 (September 17, 2004) to JD 2454258.94 (June 7, 2007). These observations were all processed with the built-in HARPS Data Reduction Software (DRS), which performs the standard bias, flat fielding, and wavelength calibration of the high-resolution spectra. The DRS also yields the extracted spectra, the cross-correlation functions (CCF), a measurement of the full-width at half maximum (FWHM), and the bisector inverse slope (BIS) (Mayor et al., 2003; D´ıaz et al., 2018). After the data passes through the DRS, radial velocity measurements are then re-computed using the HARPS-TERRA (Template-Enhanced Radial velocity Re-analysis Application) package (Anglada-Escude & Butler, 2012). Using all available orders of spectra of a star, HARPS-TERRA shifts them to zero, then co-adds the spectra, yielding one high S/N spectrum of the target. Next, the individual spectra are cross-correlated order by order against the high S/N stellar template to re-calculate the RVs. We refer to these data as TERRA hereafter.
2.2.2 TERRA2
On June 3rd, 2015, a new fiber link was installed in HARPS in order to increase the stability of the instrument and the radial velocity precision (Lo Curto et al., 2015). As a result, we treat the observations from HARPS post-upgrade as a distinct instrument and refer to these radial velocity measurements as TERRA2, since they were processed in a similar manner to the pre-upgrade HARPS data. After the upgrade, we have a total of 106 observations spanning from JD 2457559.81 (June 20, 2016) to JD 2458755.48 (September 28, 2019).
2.3 6.5m Carnegie Planet Finder Spectograph (PFS)
PFS is a high resolution, optical echelle spectrograph, which is outfitted with an $\text{I}_{2}$ cell. This Iodine absorption cell aids with wavelength calibrations to ensure precise radial velocity measurements are derived from the Doppler shift of certain spectral lines (Butler et al., 1996). PFS is temperature controlled at $27\pm 0.01^{\circ}$C and operated at ambient atmospheric pressure. The data taken prior to 2018 has a resolution of $\sim 80,000$ (0.50.5" slit). In early 2018, the old 4K x 4K CCD with 15 micron pixels was replaced with a modernized 10K x 10K chip with 9 micron pixels. Since the upgrade [^3] https://users.obs.carnegiescience.edu/crane/pfs/specs.php, observations have been taken with a 0.3” slit, yielding a resolution of $\sim 130,000$ with a radial velocity uncertainty on the order of $\sim 1$ m/s (Crane et al., 2006, 2008, 2010), for targets brighter than V=10. In this study, we have 45 observations from JD 2455847.60 (October 13, 2011) until JD 2459508.64 (October 21, 2021).
3 Stellar Properties
Table 1: Stellar Parameters of HD208487
Parameter Value Source
R.A. (J2000) $21:57:19.85$ SIMBAD 444
Dec. (J2000) 37:45:49.05-37:45:49.05 SIMBAD
$m_{V}$ 7.47 SIMBAD
BVB-V 0.82 SIMBAD
Distance (pc) $44.88$ GAIA 555
Spectral type G2V Tinney et al. (2005)
Mass $\left(M_{\odot}\right)$ $1.15\pm 0.04$ This work (ARIADNE)
Radius $\left(R_{\odot}\right)$ $1.16\pm 0.02$ This work (ARIADNE)
[…]
Luminosity $\left(L_{\odot}\right)$ $1.71\pm 0.09$ This work (ARIADNE)
[…]
$\text{Fe}/\text{H}]$ $0.09\pm 0.04$ This work (ARIADNE)
[…]
$v\sin i\left(\mathrm{km}\mathrm{s}^{-1}\right)$ $5.58\pm 0.27$ This work (SPECIES)
$\log R_{HK}^{\prime}$ -4.90 Henry et al. (1996)
44footnotetext: http://simbad.u-strasbg.fr/simbad/55footnotetext: This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.
We derived the host star properties using two independent methods based on equivalent width measurements implemented in SPECIES [^6] https://github.com/msotov/SPECIES (Soto & Jenkins, 2018) and spectral energy distribution (SED) fitting with ARIADNE [^7] https://github.com/jvines/astroARIADNE (Vines and Jenkins 2022).
3.1 SPECIES – Spectroscopic Parameters and atmosphEric ChemIstriEs of Stars
[…]
3.2 ARIADNE – spectrAl eneRgy dIstribution bAyesian moDel averagiNg fittEr
Adjusting the SED to a stellar atmosphere model is a common method to derive stellar properties. However, the physical assumptions of the underlying atmosphere model may contribute to bias the stellar parameters. Here we used a Bayesian Model Averaging approach implemented in ARIADNE (Vines & Jenkins, 2022), which fits archival photometry to several stellar atmosphere models, in order to mitigate the bias from a single atmosphere model. The stellar parameters were the results of an averaged posterior distributions from the Phoenix V2 (Husser et al., 2013), BT-Settl (Hauschildt et al., 1999; Allard et al., 2012; Kurucz, 1993; Castelli & Kurucz, 2003) SED models weighted by their respective Bayesian evidence estimates.
To derive the parameters from Table 1, SPECIES Teff, log g, [Fe/H] were used as priors along with archival photometric data, which was automatically fetched by ARIADNE for the SED fitting. ARIADNE outputs are used to derive the stellar age, mass and the equal evolutionary points from the isochrone package.
4 Periodogram Analysis
To begin the analysis of the RV data, we utilized the EXOSTRIKER package, which is a ”transit and radial velocity interactive fitting tool for orbital analysis and N-body simulations” built as an accessible GUI (Trifonov, 2019). We searched for embedded periodic signals in the data using the Generalized Lomb-Scargle (GLS) periodogram with a minimum period of 1 day and a maximum period of 10,000 days (Lomb, 1976; Scargle, 1982; Zechmeister & Kürster, 2009).
4.1 Full RV Data Set
After inserting the combined RV data from UCLES, HARPS, and PFS into EXOSTRIKER, we mean-subtracted the RVs and then fit a linear trend to the data. The top panel in Figure 2 shows the results of the original GLS periodogram, after the initial mean-subtraction and linear trend fit (with the Window function in green). The next panel displays the GLS periodogram of the residuals after fitting the first Keplerian (k1) signal, where the result shows the remaining embedded signals. Similarly, the third panel illustrates the GLS periodogram on the residuals of the combined fit of the k1 and k2 signals. Additionally, in that panel, we see flanking peaks that surround the $\sim 930$d signal. Although they appear to be statistically significant peaks, they are artifacts of the main signal due the discrete sampling of the data (window function) beating with the main signal. The bottom panel reveals that there are no remaining significant peaks after fully fitting a three-planet model.
After fitting all the signals that surpass the 0.1% FAP threshold, we find periods at $P_{1}=129.3$d, $P_{2}=493.2$d, and $P_{3}=939.7$d. This preliminary analysis helps inform what signals may appear during the more rigorous Bayesian analysis.
Figure 2: Top Panel: Original GLS periodogram of the combined and mean-subtracted RV data (black). The periodogram of the window function (sampling) for the combined data (green) is also shown. Second Panel: GLS periodogram after fitting and removing the first detected signal. Third Panel: GLS periodogram after fitting and removing both the first and second signals. Bottom Panel: GLS periodogram after fitting and removing all three previously detected signals. The blue, dotted horizontal lines represents the 0.1% False Alarm Probability level estimated from 5000 bootstrap resamplings.
4.2 Split RV Data Set
As a first check to test how stellar activity is adding noise to the RV data, or indeed causing the signals we are seeing, we repeated the periodogram analysis with split data sets, where the split is made after calculating the median S-index value in each data set (excluding UCLES data, which do not have S-index values recorded). We then split the RVs in half: one data set with all RVs corresponding to S-index values less than the median and one data set with all RVs corresponding to S-index values greater than or equal to the median. We found that the low S-index data finds all three signals, while the high S-index data only finds the inner gas giant. Thus, the low S-index data provides a significantly better fit to the data when compared to the high S-index data, which would be expected if the RV signals are all independent of the stellar activity.
4.2.1 Low S-index
The low S-index data set includes 10 TERRA, 22 PFS, and 53 TERRA2 observations, for a total of 85 RVs. Duplicating the analysis from the full data set, the GLS periodogram is used to search for any periodicity in this data set of lower stellar activity. To evaluate the models that we fit, we check the Bayesian Information Criteria (BIC). In general, a $\Delta BIC\geq 5$ ($\sim$150x more probable) is considered statistically significant (Kass & Raftery, 1995; Feng et al., 2016). After mean-subtracting the RVs and fitting a linear trend, the initial k0 model yields a BIC of 5940.85. The GLS periodogram clearly shows the first two signals above the 0.1% FAP threshold, with the third slightly below, as seen in Figure 10. Fitting the first signal, with a period of $P_{1}=129.3d$, yields a BIC of 620.87 and a $\chi^{2}_{red}=4.60$. Next, we add the second signal of $P_{2}=536.6d$ to the model, which after fully fitting ends with a period of $P_{2}=526.4d$, with a BIC of 452.98 and a $\chi^{2}_{red}=2.25$. Initially, the $P_{3}$ signal falls below the 0.1% FAP threshold. However, fitting the third period found from the highest peak and refitting all of the signals ends up reducing the BIC to 444.66, where $P_{3}=937.6d$. Adding the third signal reduced the BIC by 8.32 ($\sim$4100x improvement), enabling us to conclude that the third signal still exists in this data set. After fully fitting the three-planet model, we found that $rms=2.18m/s$, $\chi^{2}=137.94$, and $\chi^{2}_{red}=2.03$, showing a good fit to the data.
4.2.2 High S-index
The high S-index data set includes 10 TERRA, 23 PFS, and 53 TERRA2 observations, for a total of 86 RVs. The RVs corresponding to the data with higher stellar activity were examined with the GLS periodogram, finding that only the first of the previously detected peaks shows up at high statistical significance (Figure 11). A BIC of 3719.79 was found after mean-subtracting and fitting a linear trend to the RVs. Once again, we attempted to manually fit the peaks from the periods found with the full data set. After adopting the first period of $P_{1}=130.3$d as the k1 signal, the best fit is found at $P_{1}=130.6d$ with a BIC of 1020.59 and a $\chi^{2}_{red}=9.67$, showing the signal is significant within the BIC test. Neither of the remaining two signals are detected from the high S-index data. Thus, we concluded that the stellar activity is actually obscuring the signals, rather than contributing to them.
5 Bayesian Analysis
Figure 3: Top to Bottom, Left to Right: Phase-folded, best-fit model for the first, second, third, and complete Keplerian model.
Figure 4: Top, Left to Right: Final posterior distribution of the period for the first, second, and third Keplerians from the best-fitting parallel-tempered MCMC run. Bottom, Left to Right: Final posterior distribution of the semi-amplitude for the first, second, and third Keplerians from the best-fitting parallel-tempered MCMC run. The numbers at the top of each figure correspond to the mean, variance, skewness and kurtosis, median, and mode, respectively. The solid line represents a Gaussian curve with the same mean and variance.
After examining the results from the periodogram analyses, we found significant evidence for two additional planetary signals in the data, along with the previously confirmed planet, HD208487b. In order to corroborate the initial results, we followed similar procedures to those employed in Tuomi et al. (2014); Jenkins & Tuomi (2014); D´ıaz et al. (2020) to model the RVs. The Markov Chain Monte Carlo (MCMC) code, Exoplanet Mcmc Parallel tEmpering Radial velOcity fitteR [^9] https://github.com/ReddTea/astroemperor/tree/master (EMPEROR  Peña & Jenkins, in preparation), ”searches for signals in radial velocity timeseries, employing Adaptive Parallel Tempering MCMC, convergence tests and Bayesian statistics, along with various noise models.” Further, the Python code provides Bayesian posterior distribution models and the parameterization for the planetary signals to describe the best-fit orbital solution to the star system. An early version of EMPEROR has already been used in a number of works to detect small signals in RV data (Wittenmyer et al., 2017; Jenkins et al., 2019a, b; Barnes et al., 2020; Jenkins et al., 2020), where the updated version employs a custom-written Adaptive Parallel Tempering MCMC engine, similar to the EMCEE package (Foreman-Mackey et al., 2013), as used in (Gregory, 2005, 2007).
We successively ran the EMPEROR code with different properties to find the optimal setup. The statistical models included a function describing a k-Keplerian planet model, a linear trend term (to account for any existing accelerations within the system), a red-noise model (consisting of a moving average model with exponential weighted smoothing), and linear correlations with the stellar activity indices. The full mathematical description of the model with all parts included is shown below:
$y_{i,ins}$ $=\gamma_{ins}+\dot{\gamma}t_{i}+f_{k}(t_{i})+\sum_{n=1}^{q_{ins}}c_{n,ins}\xi_{n,i,ins}$
$+\sum_{l=1}^{p}\phi_{ins,l}\exp\{\frac{t_{i-l}-t_{i}}{\tau_{ins}}\}r_{i-l,ins}+\epsilon_{i,ins},$ (1)
where $y_{i,ins}$ corresponds to the observation at time $t_{i}$ for each of the distinct instruments, $\gamma_{ins}$ is the velocity offset for each of the instruments, $\dot{\gamma}$ is a linear trend term (for any existing acceleration), and $c_{n,ins}$ describes the linear correlations with $q$ stellar activity indices $\xi_{n,i,ins}$ for each instrument. $r_{i,ins}$ denotes the residuals once the model has been subtracted from the measurement. The function $f_{k}$ is a superposition of $k$-Keplerian signals denoted by:
$f_{k}\left(t_{i}\right)=\sum_{m=1}^{k}K_{m}\left[\cos\left(\omega_{m}+\nu_{m}\left(t_{i}\right)\right)+e_{m}\cos\left(\omega_{m}\right)\right]$ (2)
where $K_{m}$ is the velocity semi-amplitude, $\omega_{m}$ is the longitude of periastron, $\nu_{m}$ is the true anomaly and $e_{m}$ is the eccentricity. Since $\nu_{m}$ is a function of the orbital period and the mean anomaly $M_{0,m}$ at epoch, we can fully describe $f_{k}$ by $K_{m}$, $\omega_{m}$, $e_{m}$, $M_{0,m}$, $P_{m}$; $m\in\{1,...,k\}$.
The white noise term is denoted by the additive random variable $\epsilon_{i,ins}$, shown by:
$\epsilon_{i,ins}\sim\mathcal{N}(0,\sigma_{i}^{2}+\sigma_{ins}^{2})$ (3)
where $\sigma_{i}$ is the uncertainty reported with the measurement $y_{i,ins}$ and $\sigma_{ins}$ is the excess white noise or jitter modelled after each instrument. This nuisance parameter, which is required to model the mechanisms that create the data, has a Normal prior bound (Table 2), in the model. The remaining terms define the rest of the noise model, as described in D´ıaz et al. (2018).
5.1 Posterior Samplings and Signal Detection
Table 2: Prior Selection for the Parameters
Parameter Prior Type Range
Semi-amplitude Uniform $K\in\left(0,100\right]$
Period Uniform $P\in\left[1,8475\right]$
Eccentricity $\mathcal{N}\sim\left(0,\sigma_{e}\right)$ $e\in[0,1)$
Long. of Peric. Uniform $\omega\in[0,2\pi]$
Mean Anomaly Uniform $M_{0}\in[0,2\pi]$
Jitter $\mathcal{N}\sim\left(5,5\right)$ $\sigma_{ins}\in\left(0,100\right]$
Offset Uniform $\gamma_{ins}\in[-100,100]$
Acceleration Uniform $a\in[-1,1]$
Stellar Activity Uniform $SA\in[-1,1]$
The Adaptive Parallel Tempering provides a ladder of replicas of the system, where each step has the likelihood at a different temperature. The hotter the chain, the flatter the posterior, allowing the sampler to jump from one local maxima to another, whilst the cold chains explore the phase space in detail. Swapping states for random members of different systems makes the cold chain fully characterize each of the maxima. The adaptive scheme utilizes varying temperatures to serve as a ’good ladder.’ This ladder allows for the calculation of the Bayesian evidence via thermodynamic integration as well (more details in Vousden et al. (2016)). The prior distributions for the orbital and instrumental parameters for our model are listed in Table 2.
5.2 Model Selection
The final piece of the EMPEROR code deals with choosing the proper model. Ideally, the final model represents the most accurate description of the orbital and planetary parameters. As a result, EMPEROR performs model selection at each stage to ensure the current model actually produces an improvement over the previous model. In order for the new k+1 model to be preferred over the existing k model, we require an improvement over a statistic of choice. In this scenario, for the model comparison, the log of the evidence (Z) was chosen, and a Bayes factor of at least 5 was required, which represents a $150\times$ more probable solution than the previous model. The model evidence, or marginalized likelihood, corresponds to the likelihood of generating the data from the prior, given a correct model. In other words, it could be understood as the probability of the model itself, hence its use for model comparison (Robert, 2007). The ratio of evidences for two competing models is called the Bayes factor and is used for model comparison.
The EMPEROR MCMC code employs a burn-in stage, in which the chains search for the high-probability zones, whilst adapting their temperatures to have an efficient swap rate. When the burn-in stage is over, the samples are discarded, and temperatures stop adapting. This ensures that the chain remains ergodic whilst satisfying detailed balance. This process lasts for 80% of the total iterations. In our analysis, the Markov-chains comprised 12 temperatures, 1200 walkers, and 20000 steps, combining for 288 million iterations per model (with a 230 million iteration burn-in). Then, the chain must be longer than 50 times the autocorrelation time, or it is extended. The models we ran, beyond the straight Keplerian models include:
  1. 1.
    a white-noise model (WNO, as shown in eq. 3)
  2. 2.
    correlated noise, modelled by an exponentially weighted moving average (MAO, see last term of eq.5)
  3. 3.
    white noise with linear stellar activity correlations (CorrO, see the first sum of eq.5)
  4. 4.
    correlated noise plus the linear stellar activity correlations (CorrMA)
Initially, the Corr models were run with every available stellar index, however only the indices with correlations with the RVs that were inconsistent with zero (no matter how low the rank) were included in the final set of runs. As a result, the Corr models are based on the S-index from PFS and the BIS from TERRA1 and TERRA2.
Additionally, we ran each of these models with a fixed eccentricity of zero, a tightly constrained eccentricity described by a normal distribution centered at zero with a standard deviation of 0.1, and a constrained eccentricity described by a normal distribution centered at zero with a standard deviation of 0.3 (as explained in Jenkins et al. 2025, submitted), which enables EMPEROR to explore a wide variety of solutions based on the potential planetary eccentricities. In general, we found that fixing the eccentricity prior to zero would cause additional signals to appear that were harmonics of the real signals, primarily a consequence of the fact that the largest signal in the data corresponds to a massive planet on a non-circular orbit, and hence is poorly explained by a strict cosine-like function. All circular models that caused the 1st harmonic of this primary signal (HD208487b) to appear as a secondary signal in the data performed statistically much worse, enabling us to securely disregard these harmonic models. On the other hand, most of the models with the constrained eccentricity prior or the tightly constrained eccentricity prior yield very similar results. The main difference between the constrained and tightly constrained models are the predicted eccentricities of the outer two planets. The periods and semi-amplitudes are similar, but the constrained models tend to terminate with higher eccentricities, which results in both worse final statistics and long-term unstable planetary dynamics. In the theoretical limit of infinite walkers and steps, we would expect all of the models to converge to the same parameter configuration, even though we extended our analyses as far as reasonably possible to 288 million iterations.
We found that the CorrO with the tightly constrained eccentricity model performed the best, yielding a dynamically stable configuration. An independent verification method was performed to validate our results, which was in agreement with our best-fit model (see §5.3). The final orbital and instrumental parameters for the best-fitting model (CorrO-tightly constrained eccentricity) are shown in Table 3. The best-fit Keplerians are shown in Figure 3 and the posterior distributions for the periods and semi-amplitudes are shown in Figure 4.
5.2.1 Instrument Contributions to the Model
Due to the varying number of RVs in each dataset as well as the differences in quality of each instrument, it is worthwhile to determine each dataset’s overall contribution to the final model based on each instrument’s RMS. Each dataset’s contribution is approximated by dividing its own number of observations by the square of its RMS, ($\frac{N_{i}}{RMS_{i}^{2}}$) and then finding the weighted sum. As shown in Table 4, although AAT and PFS comprise around one fifth of the overall number of RVs, they both contribute a disproportionately low amount to the overall best-fit model. In fact, the AAT data contributes $7\times$ less value than expected based on the number of RVs. On the other side, the TERRA2 data contributes most to the model due to its high precision, and large data volume. However on the other hand, including the AAT data increases the duration of the baseline from about three to five full periods of the outer planet, which lowers the uncertainties of the orbital parameters, especially for the longer-period planets. Therefore, we find an interesting balance between modeling long-baseline, yet lower precision RV datasets, against short-baseline but high-precision measurements, whereby even though some datasets can be an order of magnitude lower in precision, their overall observing baselines can overcome these short comings when targeting longer-period worlds.
5.2.2 Comparison of Earlier Data and Later Data
[…]
Table 3: Comparison of the best WNO models (with $1\sigma$ errors) for different dataset aggregations. The overall best fit of orbital and instrumental parameters solutions for HD208487 is listed under ALL.
Parameter TERRA+TERRA2 No AAT AAT+TERRA PFS+TERRA2 AAT only ALL
$P$ (days) $129.33^{+0.04}_{-0.04}$ $129.32^{+0.03}_{-0.01}$ $129.39^{+0.15}_{-0.22}$ $129.34^{+0.01}_{-0.03}$ $129.27^{+0.23}_{-0.14}$ $\mathbf{129.36^{+0.02}_{-0.02}}$
$K$ (m/s) $18.38^{+0.08}_{-0.23}$ $18.76^{+0.27}_{-0.35}$ $17.35^{+0.95}_{-1.22}$ $18.53^{+0.10}_{-0.17}$ $16.70^{+0.95}_{-0.02}$ $\mathbf{18.28^{+0.29}_{-0.40}}$
$e$ $0.36^{+0.01}_{-0.02}$ $0.35^{+0.01}_{-0.01}$ $0.18^{+0.04}_{-0.06}$ $0.35^{+0.00}_{-0.01}$ $0.13^{+0.03}_{-0.02}$ $\mathbf{0.37^{+0.01}_{-0.01}}$
$\omega$ (rad) $1.21^{+0.03}_{-0.01}$ $1.21^{+0.01}_{-0.03}$ $41^{+0.55}_{-0.35} 1.1$$16^{+0.06}_{-0.04}$ $2.58^{+0.48}_{-0.65}$ $\mathbf{1.10^{+0.01}_{-0.01}}$
$M_{0}$ (rad) $3.43^{+0.08}_{-0.10}$ $3.40^{+0.07}_{-0.01}$ $1.83^{+0.38}_{-0.68}$ $3.23^{+0.04}_{-0.05}$ $0.58^{+0.78}_{-0.49}$ $\mathbf{1.95^{+0.05}_{-0.04}}$
$m\sin i$ ($M_{J}$) $0.47^{+0.01}_{-0.01}$ $0.48^{+0.01}_{-0.01}$ $0.47^{+0.03}_{-0.03}$ $0.47^{+0.01}_{-0.01}$ $0.45^{+0.03}_{-0.00}$ $\mathbf{0.46^{+0.01}_{-0.00}}$
$AU1$ $0.52^{+0.00}_{-0.00}$ $0.52^{+0.00}_{-0.00}$ $0.53^{+0.00}_{-0.00}$ $0.53^{+0.00}_{-0.00}$ $0.52^{+0.00}_{-0.00}$ $\mathbf{0.53^{+0.00}_{-0.00}}$
$P_{2}$ (days) $490.82^{+5.62}_{-0.31}$ $488.10^{+0.89}_{-1.54}$ $934.56^{+17.97}_{-42.29}$ $485.60^{+4.71}_{-6.68}$ $923.92^{+26.38}_{-19.30}$ $\mathbf{923.06^{+2.02}_{-2.76}}$
$K_{2}$ (m/s) $3.73^{+0.27}_{-0.01}$ $4.35^{+0.33}_{-0.46}$ $9.76^{+1.26}_{-1.98}$ $4.03^{+0.17}_{-0.09}$ $8.69^{+1.13}_{-0.21}$ $\mathbf{6.18^{+0.09}_{-0.07}}$
$e_{2}$ $0.17^{+0.02}_{-0.06}$ $0.30^{+0.06}_{-0.10}$ $0.17^{+0.06}_{-0.12}$ $22^{+0.01}_{-0.04} 0.0$$08^{+0.03}_{-0.03}$ $\mathbf{0.19^{+0.01}_{-0.02}}$
$\omega_{2}$ (rad) $4.55^{+0.01}_{-0.40}$ $4.72^{+0.33}_{-0.48}$ $3.89^{+1.20}_{-0.04}$ $4.00^{+0.14}_{-0.17}$ $2.94^{+1.46}_{-1.10}$ $\mathbf{5.96^{+0.30}_{-5.96}}$
$M_{02}$ (rad) $0.74^{+0.90}_{-0.98}$ $0.41^{+0.19}_{-0.14}$ $3.61^{+1.34}_{-1.70}$ $2.39^{+0.00}_{-0.32}$ $3.96^{+0.00}_{-1.07}$ $\mathbf{1.29^{+0.52}_{-0.29}}$
$m_{2}\sin i$ ($M_{J}$) $0.16^{+0.01}_{-0.01}$ $0.18^{+0.01}_{-0.01}$ $0.51^{+0.07}_{-0.07}$ $0.17^{+0.01}_{-0.01}$ $0.46^{+0.06}_{-0.01}$ $\mathbf{0.32^{+0.01}_{-0.01}}$
$AU2$ $1.28^{+0.01}_{-0.01}$ $1.27^{+0.01}_{-0.01}$ $1.96^{+0.03}_{-0.03}$ $1.27^{+0.01}_{-0.01}$ $1.95^{+0.04}_{-0.03}$ $\mathbf{1.94^{+0.01}_{-0.01}}$
$P_{3}$ (days) $959.17^{+419.38}_{-194.63}$ $941.38^{+28.27}_{-18.51}$ $1294.65^{+1191.79}_{-38.62}$ $979.74^{+33.79}_{-36.32}$ $186.57^{+1125.53}_{-186.57}$ $\mathbf{1380.13^{+19.20}_{-8.25}}$
$K_{3}$ (m/s) $3.16^{+1.01}_{-0.94}$ $3.54^{+0.36}_{-0.50}$ $5.61^{+1.63}_{-3.77}$ $3.09^{+0.42}_{-0.37}$ $4.35^{+0.07}_{-1.57}$ $\mathbf{2.46^{+0.30}_{-0.09}}$
[…]
$\omega_{3}$ (rad) $1.08^{+2.25}_{-1.69}$ $0.57^{+2.24}_{-1.65}$ $2.80^{+1.49}_{-0.65}$ $3.34^{+0.12}_{-1.12}$ $4.91^{+1.86}_{-3.09}$ $\mathbf{0.29^{+0.28}_{-0.09}}$
$M_{03}$ (rad) $3.30^{+0.80}_{-1.37}$ $3.27^{+0.37}_{-0.82}$ $5.47^{+1.85}_{-3.32}$ $5.67^{+2.17}_{-4.10}$ $3.79^{+0.37}_{-2.08}$ $\mathbf{2.61^{+0.47}_{-0.42}}$
$m_{3}\sin i$ ($M_{J}$) $0.168^{+0.07}_{-0.07}$ $0.18^{+0.02}_{-0.02}$ $0.33^{+0.10}_{-0.10}$ $0.16^{+0.02}_{-0.02}$ $1.34^{+0.11}_{-0.10}$ $\mathbf{0.15^{+0.01}_{-0.02}}$
$AU3$ $1.995^{+0.25}_{-0.25}$ $1.97^{+0.03}_{-0.03}$ $2.436^{+1.52}_{-1.52}$ $2.00^{+0.05}_{-0.05}$ $0.67^{+1.78}_{-0.67}$ $\mathbf{2.54^{+0.02}_{-0.01}}$
$\gamma_{UCLES}$ (m/s) - - $18.06^{+2.71}_{-1.83}$ - $21.28^{+0.06}_{-1.53}$ $\mathbf{1.78^{+0.37}_{-0.94}}$
$\gamma_{TERRA1}$ (m/s) $1.18^{+2.07}_{-0.19}$ $1.87^{+0.69}_{-0.16}$ $31.28^{+1.79}_{-0.21}$ - - $\mathbf{12.10^{+0.41}_{-0.76}}$
$\gamma_{TERRA2}$ (m/s) $18.20^{+28.31}_{-15.94}$ $14.78^{+2.93}_{-2.50} -$ $3.91^{+0.20}_{-0.91}$ - $\mathbf{-8.373^{+0.74}_{-1.99}}$
$\gamma_{PFS}$ (m/s) - $15.86^{+3.25}_{-2.49} -$ $3.18^{+0.30}_{-0.88}$ - $\mathbf{-8.65^{+0.70}_{-2.41}}$
[…]
$\sigma_{UCLES}$ (m/s) - - $94^{+0.99}_{-0.70} - 4.4$$41^{+0.92}_{-0.63}$ $\mathbf{5.65^{+0.81}_{-0.67}}$
$\sigma_{TERRA1}$ (m/s) $2.03^{+0.45}_{-0.38}$ $1.93^{+0.50}_{-0.37}$ $2.49^{+0.03}_{-0.36}$ - - $\mathbf{1.61^{+0.18}_{-0.13}}$
$\sigma_{TERRA2}$ (m/s) $33^{+0.09}_{-0.05} 1.1$$13^{+0.24}_{-0.14} -$ $1.25^{+0.14}_{-0.13}$ - $\mathbf{1.38^{+0.17}_{-0.15}}$
$\sigma_{PFS}$ (m/s) - $3.15^{+0.58}_{-0.45}$ - $3.66^{+0.10}_{-0.31}$ - $\mathbf{3.12^{+0.29}_{-0.12}}$
BIC 595.540 850.161 469.948 758.665 358.478 1209.520
$\log{Z}$ -288.352 -424.660 -223.560 -369.246 -169.400 -582.577
RMS 1.806 2.431 4.568 2.553 4.915 3.741
$\chi^{2}_{red}$ 1.117 1.203 1.188 1.103 1.467 1.172
Table 4: Contribution of each Instrument to the Best-fit Model.
AAT TERRA PFS TERRA2
RVs 48 20 45 106
RMS 6.57 1.62 3.58 1.85
RV dataset % 22 9 21 48
Contribution % 3 18 8 72
5.2.3 AAT-only Run
Ensuring consistency between earlier research and the current analysis remains an essential aspect of follow-up studies. Initially, Tinney et al. (2005) detected HD208487b utilizing a Periodogram analysis. Later, Gregory (2005, 2007) claimed the detection of HD208487c with a more statistically robust MCMC method compared to Tinney’s Periodogram analysis. To maintain consistency, we re-analyzed the AAT data alone with EMPEROR. The results from these MCMC runs, summarized in Table 3, indicate very similar findings, adding confidence to the past research.
5.3 Independent verification of results
In order to increase the robustness and reliability of the results, we analyzed the data with independent sampling methods and codes. We applied the delayed-rejection adaptive Metropolis algorithm (Haario et al., 2006) to detect maxima in the posterior probability density and the simple adaptive-Metropolis algorithm (Haario et al., 2001) to estimate parameters and obtain solutions (see e.g. Butler et al., 2017; Tuomi et al., 2018).
We also increased the set of available activity-indices by calculating the ”differential velocities” of Feng et al. (2017) for the TERRA reduction of HARPS. This is to say that we divided the 72 HARPS orders into three equal subsets, corresponding to the red-most, middle, and blue-most parts of the corresponding spectra. These were then used independently to obtain three sets of radial velocities denoted as $y_{i,1},y_{i,2}$, and $y_{i,3}$ for $i=1,...,N_{\rm HARPS}$. The differential velocities are thus $\xi_{i,1}=y_{i,1}-y_{i,2}$ and $\xi_{i,2}=y_{i,2}-y_{i,3}$ that can be used as additional activity indices. Because any and all Doppler signals of planets have been subtracted when producing these indices, the remaining variability simply corresponds to noise and activity-induced variations. In particular, wavelength-dependent variations that are present in these indices enable one to filter them out from the data by accounting for the respective correlations with these indices in accordance with Eq. (5). The models are thus equivalent except for the fact that the additional activity indices enable removing some additional activity-induced variability. Such variability indeed was found to be only weakly present in the HARPS data. Yet, since the results were consistent, the postulated planetary signals presented in the manuscript show no evidence for variations as a function of spectral wavelength.
[…]
6 Stellar Activity and Correlations
Following a similar strategy as in Santos et al. (2014) and D´ıaz et al. (2018), we investigated how stellar indices might be affecting or contributing to the periodic signals found within the RVs. First, we checked the GLS periodograms of every available stellar index to see if any stellar activity peak is associated with a previously found periodic signal, or its harmonics, as shown in Figure 5. No signals are significant above the 0.1% FAP threshold for any index in TERRA or PFS, as well as the Bisector Inverse Slope (BIS) of TERRA2. The S-index in TERRA2 has significant peaks in the mid 20 days and also around 800 days. Also, the Full Width at Half Maximum (FWHM) from TERRA2 has a significant peak at $\sim 22$d. However, none of these peaks approach the previously found periodic signals. The Combined S-index and Combined FWHM both had significant peaks at long periods around $\sim 4000$d. Additionally, the RV vs stellar index correlations were plotted in Figure 6. Table 5 shows the calculated Pearson rank and p-value of the RV/index correlations. Generally, in order for a correlation to be statistically significant, the p-value must be $p\leq 0.05$. We note that the TERRA S-index, TERRA2 S-index, and TERRA2 FWHM have p-values approaching significance, however none have a rank corresponding to even a moderate correlation. A rank of $\lvert r\rvert\geq 0.5$ is required to be considered a moderate or stronger correlation. Since none of the stellar index correlations have $p\leq 0.05$ nor $\lvert r\rvert\geq 0.5$, there are no significant nor strong/moderate correlations between the RVs and the stellar indices.
Table 5: Stellar Activity Index Correlations of HD208487
Instrument Index Pearson rank p-value
TERRA S-index 0.365 0.113
TERRA FWHM 0.207 0.381
TERRA BIS 0.163 0.492
PFS S-index 0.209 0.173
TERRA2 S-index 0.189 0.052
TERRA2 FWHM 0.156 0.110
TERRA2 BIS 0.122 0.212
Figure 5: GLS periodogram of the available stellar activity indices. From top to bottom: PFS S-index, TERRA S-index, TERRA FWHM, TERRA BIS, TERRA2 S-index, TERRA2 FWHM, TERRA2 BIS, Combined S-index, Combined BIS, Combined FWHM. The green, solid vertical lines show the best-fit periods of the three planetary signals. The blue, dotted horizontal lines represents the 0.1% False Alarm Probability (FAP) level estimated from 5000 bootsrap resamplings.
Figure 6: Correlations of the available stellar activity indices. From top to bottom, left to right: PFS S-index, TERRA S-index, TERRA FWHM, TERRA BIS, TERRA2 S-index, TERRA2 FWHM, TERRA2 BIS. The lines show the best-fit correlation between the activity indicator and the RVs.
6.1 Fitting Stellar Activity with EMPEROR
Table 6: EMPEROR results for Combined FWHM for HD208487 with $1\sigma$ errors.
Activity Index: Combined FWHM (ecc 0.3 prior)
k1 k2
[…]
BICBIC -892.99 -902.66
To expand our search for periodic RV modulations induced by stellar activity, we incorporated Bayesian statistical analysis. As a novel approach, we executed the EMPEROR code on the available stellar activity indices (the authors are unaware of any previous procedures involving Bayesian statistical analysis on stellar activity indices). Utilizing the EMPEROR code enables a more rigorous analysis of the stellar activity signals compared with the standard GLS and correlation analysis. Moreover, the posterior distributions derived from the Markov chains afford a comprehensive assessment of certainty across the entire period space, facilitating a thorough evaluation of the distinctness of the signals.
We completed several EMPEROR runs for each stellar activity index. Each index was initially analyzed individually, before joining the datasets of similar indices from different instruments into one run to test the stability and significance of the signals over the full baseline of the activity data. Further, we tested the differences between how EMPEROR fits the stellar activity indices with the fixed zero eccentricity and the constrained ($\mathcal{N}\sim(0,0.3,0,1$)) eccentricity prior. The results of the noteworthy EMPEROR run is summarized in Table 6.
[…]
7 Photometric Analysis
7.1 ASAS Photometry
In order to provide additional analysis of this star/planetary system, All-Sky Automated Survey (ASAS) photometric data were analysed (Pojmanski, 1997). Searching through the ASAS archive, we found 608 V-band observations spanning just over nine years, between JD 2451870.55 (November 22, 2000) and JD 2455168.58 (December 3, 2009). Following the ASAS guidelines, only data marked with an ”A” or ”B” quality flagged are kept (henceforth all data), leaving 509 usable data points. However, when examining the photometric time series, we see evidence of instrumental issues in the first $\sim 300$d of observations, see Figure 7 top. In order to clean the data, we calculated the mean and standard deviation of all data after JD 2452300. Then all data falling within three standard deviations of the calculated mean for the entire baseline were kept, as seen in Figure 7 middle. Finally, the GLS periodogram was calculated from the cleaned data, where we sampled the period space from 1 day to 10,000 days, performing 80,000 period samples.
Table 7: The top five peaks in the cleaned ASAS photometry data with their corresponding FAP.
Period FAP
29.50 0.21
28.75 0.39
27.25 0.52
330.60 0.54
384.47 0.84
[…]
Figure 7: Left (Right): ASAS V-band photometry (Hipparcos Photometry). Top: ”A” and ”B” flagged V-band ASAS photometry (all Hipparcos data) for HD208487. Middle: Cleaned V-band ASAS photometry (Hipparcos) for HD208487. Bottom Left: Generalized Lomb-Scargle periodogram of the photometry. Horizontal lines correspond to the 0.1% (dotted blue) FAP significance level computed via 5000 bootstrap iterations on the photometry time series. The vertical green lines represent the periods of the three planets.
Figure 8: GLS periodogram of cleaned ASAS Photometry. Top: Original GLS periodogram of combined, mean-subtracted ASAS Photometry data. Second: GLS periodogram after fully fitting the first signal. Third: GLS periodogram after fully fitting the second signal. The blue, dotted horizontal lines represents the 0.1% FAP level estimated from 5000 bootsrap resamplings.
7.1.1 Fitting Stellar Cycles
Extending our analysis of the ASAS photometry, we searched for periodic stellar cycles in the cleaned data. After noting that none of the GLS periodogram peaks align with the planetary signals, we fit the strongest periods, as shown in Figure 8. The segment of the periodogram with the highest power is continuously increasing, with no defined peak, even as it approaches a period of 10,000 days. Thus, we conclude that the signal is likely a baseline effect, rather than a real cycle. The next two peaks constitute very similar periods of 29.51 days and 28.73 days. Based on the strong peak in ASAS and a very similar period in the GLS periodogram of the S-index of TERRA2, we conclude that the spikes in this period range are very likely to be due to the stellar rotation period, possibly even being affected by differential rotation. With continued spectral monitoring of this star we can further test this hypothesis in the future.
7.2 Hipparcos Photometry
To further explore the photometric properties of HD208487, we investigated the publicly available Hipparcos data (Hip, 1997). These data were obtained from the NASA Exoplanet Archive [^10] https://exoplanetarchive.ipac.caltech.edu/, which contained 90 data points, spanning three years from JD 2447881.99 (December 21, 1989) to JD 2448977.85 (December 21, 1992). Following the data quality flag information, 1 point was identified as low-quality and unusable data. After removing the poor data point, the GLS Periodogram was calculated, in which we sampled the period space from 1 day to 10,000 days, performing 80,000 period samples. There were no peaks coming close to statistical significance, demonstrating a lack of any periodic photometric signals in the Hipparcos data.
8 Discussion
With 184 new RV observations spanning 15+ years since the previous analyses (Tinney et al., 2005; Gregory, 2005; Goździewski & Migaszewski, 2006; Gregory, 2007; Wright et al., 2007; Babu et al., 2010), this system is ripe to study and to re-evaluate the reality of the planetary system. Using the SPECIES and ARIADNE codes, we provide updated stellar parameters to HD208487. In order to fully understand the RV dataset and best characterize the system, we utilized GLS periodograms and EMPEROR parallel-tempering MCMC samplings, which revealed new periodic signals embedded in the RVs. An independent, analysis accounting for wavelength-dependent variations in the HARPS-TERRA velocities yielded consistent results providing confidence in the final parameters. The best-fitting model to the current RVs reveals a candidate three-planet system.
Our best-fit model, which confirms the existence of HD208487c, alters some of the parameters. Although the period, $923.06^{+2.02}_{-2.76}$d, falls within the uncertainty published by Gregory (2005, 2007), the semi-amplitude, $6.18^{+0.09}_{-0.07}$m/s, and mass, $0.32\pm 0.01m_{j}$, are now slightly lower leading to a classification as a Saturn-mass planet. Further, our analysis reveals a third signal, relating to a candidate outer planet, HD208487d, with a period of $1380.13^{+19.20}_{-8.25}$d, a semi-amplitude of $2.46^{+0.30}_{-0.09}$m/s, and a minimum mass of $0.15^{+0.01}_{-0.02}m_{j}$. The two outer signals are found very close to the 3:2 period ratio ($P_{d}/P_{c}=1.495$), further increasing the reliability of these as being due to Doppler signals, since it is difficult to produce 3:2 frequency ratio signals through aliasing (e.g. Hara et al., 2020), particularly with a possible 2:1 signal also present in the data, and also transit observations have shown such period ratios to be fairly common (Lissauer et al., 2011; Fabrycky et al., 2014; Lissauer et al., 2024). These tests therefore leave us with a system that presents an interesting and unique system architecture, as there are currently no other known planetary systems that contain two planets on 500-1500d orbits that are between $0.1-0.4m_{j}$. HD208487 contains an inner gas giant along with a resonant pair of outer, Saturn/sub-Saturn planets that are close to a second order period resonance.
As part of our analysis, we examined all of the available stellar activity indices, including, to our knowledge, the first use of Bayesian analysis on stellar activity indices to check for activity-induced RV signals. We find that none of the stellar indices have peaks matching the RV periods of the discovered planets, nor their harmonics. Although the Bayesian analysis of the stellar activity indices revealed one signal (Combined FWHM) in proximity to the RV signals, it is statistically distinct based on the current data. Additionally, none of the stellar activity indices are statistically significantly correlated with the observed RVs, nor reach the level of even a moderate correlation, which was not the case for analyses done on stars like HD26965, for example D´ıaz et al. (2018). Furthermore, during the GLS periodogram analysis with EXOSTRIKER, we split the RV data sets into two separate sets, based on their S-index values. Although only the inner gas giant was detected in the active dataset, three signals were found in the inactive dataset, the inner gas giant, the $\sim 500d$ signal, and the Saturn-mass planet, HD208487c. Since the model fit of the inactive set was significantly better, the stellar activity may actually be masking the Doppler signals rather than contributing to them. Again, we reiterate that more data is needed to confirm the nature of the intermediate, $\sim 500d$ signal.
In order to evaluate the stability of the signals over time, we split the data into early and later subsets. HD208487b remains entirely consistent in both data sets. HD208487c is also found in both data sets, with the period and semi-amplitude remaining consistent within uncertainties. The major difference between the earlier and later data sets relates to the period of the third significant signal. The earlier data yields the longest-period signal, while an intermediate-period signal arises in the later data. Both the overall best-fit EMPEROR model and an independent analysis favor the longest-period signal, leading to the adoption of HD208487d as an outer planet. However, we note that often times our Bayesian analyses arrived at a model favoring the solution with the shorter-period planet, yet even in the best case scenarios the model probabilities were far below the models containing the outer signal. Yet this might indicate that there is indeed another Doppler signal close to a 2:1 period ratio with the HD208487c signal, and from a dynamical standpoint this also may be favored, since it would be another example of a dynamically packed planetary system.
Extending our investigation of this system, we checked ASAS photometry data to search for any periodic variations in the star’s visible magnitude. Examining the data showed no significant peaks within the photometry. However, the highest power peaks, paired with stellar activity indices from TERRA2, suggest a stellar rotational period around $P_{rot}\sim 29$d.
8.1 Previous Signal Conclusions
After the discovery of HD208487b by Tinney et al. (2005) and the subsequent detection of HD208487c by Gregory (2005), some disagreements about the planetary configuration arose. Gregory (2007), Goździewski & Migaszewski (2006), and Wright et al. (2007) also detected HD208487c as a moderately eccentric ($e\simeq 0.4$), sub-Jovian world, however the latter two papers find equally probable solutions at $P_{c}\sim 14.5$d, $P_{c}\sim 28$d, or as an activity-induced RV signal, leading them to disregard the existence of HD208487c as an outer planet. Babu et al. (2010) also claim a potential planet around $P_{c}\sim 14.5$d.
Conclusions drawn from our analysis, applying updated analysis methods and much more data over a longer baseline and at high RV precision, posits that the stellar rotation period is around $P_{rot}\sim 29$d, possibly with some evidence for differential rotation, and therefore this leads to the conclusion that Goździewski & Migaszewski (2006); Wright et al. (2007); Babu et al. (2010) actually found the rotation period and a likely alias at half this rotation period. Further, we believe that the detection of HD208487c as a moderately eccentric, sub-Jovian world fits the original data, as seen in Section 5.2.3. The new, higher precision RV data collected over the last 15+ years now enables the detection of candidates HD208487d, which would be an outer super-Neptune/sub-Saturn planet, and HD208487c, which would be a Saturn-mass planet in a near-circular orbit, explaining the different planetary models discussed in these works.
8.2 Habitable Zone Planets
As we have found new planets in this work, we can try to more deeply understand and characterize them. Without observing a transit, we do not have a way to constrain their radii, meaning we can’t determine their densities or physical parameters. However, since their orbits are well-constrained, we can gain a better understanding of this multi-planet system. Based on the definition described by Kasting et al. (1993) and Kopparapu et al. (2013), we can measure the limits of the Habitable Zone (HZ) of HD208487, using the following equations:
$S_{eff}=S_{eff\odot}+aT_{\star}+bT^{2}_{\star}+cT^{3}_{\star}+dT^{4}_{\star}$ (4)
where $S_{eff}$ is the effective solar flux, $T_{\star}=T_{eff}-5780K$ ($T_{eff}$ is the stellar effective temperature), and a, b, c, d, and $S_{eff\odot}$ are constants whose values are given in Table 3 of Kopparapu et al. (2013).
$d=\sqrt{\frac{L_{\star}/L_{\odot}}{S_{eff}}}$ (5)
where $L_{\star}$ and $L_{\odot}$ is the luminosity of HD208487, in this case, and the Sun, respectively, and the radius, d, is measured in astronomical units (AU). We used the coefficient values from the ”recent Venus” and ”early Mars” models (Kopparapu et al., 2013). From the results of the ARIADNE code, the luminosity of HD208487 is $L_{\star}=1.71L_{\odot}$. As a result we found that the HZ ranges from 0.97AU to 2.25AU. We find that HD208487c would orbit the star at around 1.94AU and HD208487d would orbit at around 2.54AU, so HD208487c would lie within the HZ, while HD208487d would orbit just outside the proposed range of the HZ (Jones et al., 2006).
8.3 Exomoons
Long period exoplanets are also interesting as they could potentially host natural satellites (also known as exomoons). This is because, for the case of planets in close-in orbits, the extreme gravitational interactions with the host stars could destabilize potential exomoon orbits (e.g. Namouni, 2010; Spalding et al., 2016; Dobos et al., 2021). Moons are thought to be formed through one of two processes, i.e. coaccretion in the circumplanetary discs around the young planets (e.g. Heller et al., 2014; Barr, 2016; Nakajima et al., 2022) and planetary collisions/captures (Heller et al., 2014; Barr, 2016; Barr & Bruck Syal, 2017). Since coaccretion in the circumplanetary discs might require larger amount of materials to form moons, this process might be possible in the case of larger Jupiter-sized planets, as could be the case with HD208487c. However, for HD208487d, which appears to be a Neptunian-sized, planetary collisions/capture could potentially lead to a moon formation.
[…]
Since both HD208487c and HD208487d signals have been detected only through the RV method, we can only estimate the minimum mass, and thus, the minimum radius of the Hill sphere for them, which we have estimated to be 12.935 MKm and 13.156 MKm respectively. The minimum distance from the host-star within which a moon can survive in a prograde orbit around HD208487c and HD208487d are 4.993 MKm and 5.599 MKm respectively, and those for a retrograde orbit are 9.57 MKm and 10.786 MKm respectively. For comparison, the distance of the outermost Galilean moon, Callisto, from Jupiter is 1.883 MKm, and the largest retrograde satellite in our Solar system, Triton, orbits around Neptune only at a distance of 0.354 MKm. Thus, both HD208487c and HD208487d are capable of hosting potential exomoons, and since they are both orbiting in the HZ around their host star, potential presence of HZ exomoons make them even more interesting. However, detecting such exomoons would require extremely high precision photometric and spectroscopic follow-ups, which could only be possible using the next generation large telescopes (e.g. Simon et al., 2009; Kipping, 2009; Timmermann et al., 2020; Saha & Sengupta, 2022; Saha, 2022, 2024).
8.4 Dynamics
Figure 9: Temporal evolution of a six-planet simulation showing a merger between the first three and the fourth and fifth planets, leaving behind a 3 planet system with similar masses to the observations. The upper panel displays the semi-major axis $a_{i}$, the variation of the apocenter $Q$ (defined as $a_{i}[1+e_{i}]$), and the pericenter ($q=a_{i}[1-e_{i}]$). In the bottom panel we show the variation of the eccentricities.
Motivated by the apparent architecture of the HD208487 system, and with the aim of testing if such a system could be possible and understandable, we explore the possibility that the system’s origin could be attributed to a dynamical instability of an initially more ordered system. The star hosts two moderately eccentric inner planets ($e_{1}\sim 0.37$ and $e_{2}\sim 0.2$), each three and two times more massive than the outer Neptune-like planet. The two outer planets are near resonance, while the innermost giant appears dynamically detached from the outer two.
We posit that this more ordered system consisted of a coplanar six-planet system orbiting a $M_{\star}=1.15M_{\odot}$ central star in a near-resonant chain configuration (with consecutive period ratios near 3:2). We envision a scenario in which a six-planet system was formed in a gaseous disk and got captured in resonance due to planet-disk interactions. After dispersal of the disk, the resonance chain is broken and the system becomes unstable—a scenario reminiscent to breaking of the resonance chains applied to super-Earths in the Kepler sample (Izidoro et al., 2017). However, unlike the Kepler planets where instabilities mainly lead to mergers between planets, the cold super-Neptunes in HD208487 may lead to either collisions between planets or ejections from the system. We can quantify their relative probabilities using the so-called Safronov number for parabolic orbits as
$\theta^{2}$ $\equiv$ $\left(\frac{2GM_{p}}{R_{p}}\right)\left(\frac{a_{p}}{GM_{\star}}\right)$ (6)
$=$ $\left(\frac{M_{p}}{M_{\rm J}}\frac{R_{\rm Jup}}{R_{p}}\frac{a_{p}}{0.25{\rm au}}\right)\left(\frac{M_{\odot}}{M_{\star}}\right),$
where for $\theta^{2}\gg 1$ the planetary radius is so small that close encounters mostly lead to ejections relative to collisions, while for $\theta^{2}<1$ collisions are more frequent (Ford et al., 2001; Petrovich et al., 2014). A typical super-Earth at 0.5 au has $\theta^{2}\sim 0.25$ leading to collisions and small eccentricity excitation (Ford et al., 2001; Petrovich et al., 2014), but a cold Neptune at $\sim 2$ au has $\theta^{2}\sim 1$ leading to mixed outcomes and potential mergers with significant eccentricity excitation. We numerically evaluate this possibility next.
We perform 250 simulations using the REBOUND code (Rein & Liu, 2012) with the primary objective of investigating the possibility of achieving an orbital configuration resembling the observed data, thereby deviating from the conventional ‘peas in a pod’ scenario. The planetary masses are $m_{i}=0.15\,m_{\rm Jup}$ with $i=1\dots 6$, and we place the planets 5% away from the exact 3:2 - 3:2 - 3:2 - 3:2 - 3:2 resonant chain. We chose 5 values for the initial eccentricities for the two inner planets: $e_{1}=e_{2}=0.05,0.1,0.15,0.2,0.25$ to speed up the phase of dynamical instability, while the outer four bodies start with a fixed value of 0.05. For each of them we randomly selected 50 values for the initial argument of periastron $\omega_{i}\in[0,2.\pi]$ and inclinations randomly distributed between $0^{\circ}$ and $2^{\circ}$. Collisions happen whenever the distance between two planets is less than the sum of their radii: $d<r_{i}+r_{j}$, where the radius of each planet follows the $r_{i}\propto m_{i}^{-2.06}$ law by Lissauer et al. (2011).
Over a simulation period of $5\times 10^{5}$ years, none of the systems remained stable and maintained their six-planet configuration. In this compact configuration, all the simulations underwent an instability leading to planet collisions generally leaving behind 3, 4, and 5 planets. We note that the rate of stable (unstable) systems is highly sensitive to the initial conditions and will likely decrease (increase) as our simulations continue to evolve for longer timescales. Around $50\%$ of the systems that underwent an instability, led to the formation of a three-planet system. Of these, in $\sim 20\%$ of cases, the three inner planets merged as a massive innermost body, and the fourth and fifth planets merged as an intermediate mass middle planet, resulting in a three-planet system with an architecture similar to the one observed in our data, with inclinations reaching a final root mean square (RMS) of $1^{\circ}$. Figure 9 illustrates the temporal evolution of a representative simulation where, at $\sim 5\times 10^{4}$ years, a collision occurs leading to the merging of the inner 2 bodies into a single planet, after which a second collision occurs at $\sim 10^{5}$ years, finally forming the innermost body, $m_{1}$. Around $\sim 5\times 10^{4}$ years, a collision of the $4^{th}$ and $5^{th}$ forms the middle planet, $m_{2}$, while the outermost body never collides with other bodies. The top panel displays the semi-major axis, peri- and apocentric distances of each planet ($a_{i}(1\mp e_{i})$, respectively), while the bottom panel showcases the eccentricities as a function of time. The post-merger planet attains high eccentricities, providing a potential explanation for the observed eccentricities of the inner planets.
Although the approximation of treating collisions as perfectly inelastic mergers might not capture the complete realism of the scenario, it provides valuable insights as a proof of concept for understanding how such a system can form. As already stated by Izidoro et al. (2017), once the gas dissipates resonant chains may become dynamically unstable, undergoing a phase of gravitational scattering. Our simulations begin after the dissipation of the gas, exploiting the phase of dynamical instability, allowing the formation of a system featuring a fairly eccentric massive inner planet, largely separated from the outer pair.
9 Conclusions
This paper provides a thorough re-investigation into the HD208487 planetary system. First, we derived an updated set of stellar parameters utilizing the stellar modeling codes, SPECIES and ARIADNE. Next, we comprehensively analyzed the currently available RV data of HD208487 to search for signals not previously found in older investigations. Further, we developed a more rigorous methodology to analyze stellar activity indices. Finally, we evaluated ASAS and Hipparcos data to search for any periodic photometric signals.
As a result of our RV analysis, we determined that HD208487 possibly contains at least three planets in the system (see Table 3). We provided updated parameters to the sub-Jovian HD208487b with a period of 129.36129.36d. Formerly believed to be a possible sub-Jovian and moderately eccentric planet (Gregory, 2005), we now show HD208487c is actually a candidate Saturn with a period of 923.06923.06d and a near-circular orbit. Additionally, we discovered a new super-Neptune/sub-Saturn candidate, HD208487d, a planet that would be orbiting with a period of 1380.131380.13d. HD208487c and HD208487d have an orbital period ratio of $P_{d}/P_{c}=1.495$, indicating they are very close to a 3:2 MMR. There is also some evidence that leads us to believe that there may exist a 4th planet in this system at close to the 2:1 period ratio with HD208487c, which would further dynamically pack the planetary system, however more long-baseline RV data will be required to confirm this as a reality or not.
Our methodology to analyze the available stellar activity indices with a Bayesian statistical analysis enabled us to distinguish the RV signals from the statistically significant activity signals. We found no correlations between the RVs and stellar activity signals and no peaks in the GLS at the periods of the three planets. Using EMPEROR to provide a more accurate description, we found that the nearest detected stellar activity signal is separated by $7.1\sigma$ from $P_{c}$. Thus, we determined that there is a high probability that the RVs and stellar activity signals are distinct, based on the current data. An independent analysis of the RVs using colour-dependent terms to detrend sources of colour dependent noise also did not yield any correlations, arriving at the same system configuration as that of the EMPEROR analysis.
Concluding our observational analysis, we studied the ASAS and Hipparcos photometric data. After fitting the long-period, baseline-effect signal, we found a statistically significant period at $\sim 29$d. Combining this result with our GLS and EMPEROR analysis of the stellar activity indices, we believe this signal to indicate the rotation period of the star, with potential differential rotation. This result explains the false RV signals, and a likely half-period alias, posited by Goździewski & Migaszewski (2006), Wright et al. (2007), and Babu et al. (2010).
Finally, we performed dynamical simulations in an effort to better understand the formation history of this complex system. By assuming the common ‘peas in a pod’ configuration to begin with, including additional planets in the system close to a resonant chain, we were able to show that the inner moderately eccentric gas giant could be the remnant of a previous collision of lower-mass planets. Of the systems that underwent an instability and subsequent planetary collisions, such a scenario occurred $\sim 20\%$ of the time in our simulation set. This scenario originates from systems with Safranov numbers close to unity, that allows for significant scattering and eccentricity excitation before the merger occurs, possibly representing the breaking point of more ordered configurations and providing a chaotic look at the final architectures of planetary systems. Systems like HD208487 in our simulations, end up with relatively low inclinations ($RMS\sim 1^{\circ}$), similar to the inclinations of Kepler planets within 3-planet systems (Fabrycky et al., 2014; Zhu et al., 2018). However, unlike Kepler planets, HD208487 planets orbit with semi-major axes >1>1 au, unreachable by Kepler. Future missions like PLATO may spot systems with these architectures.
Acknowledgements.
Part of this work was supported by the United States Fulbright Fellowship in partnership with the Fulbright Chile Commission. JSJ gratefully acknowledges support by FONDECYT grant 1201371, from the ANID BASAL project FB210003, and from CASSACA grant CCJRF2205. CC was supported by FONDECYT grant n 3230283. CP acknowledges support from CATA-Basal AFB-170002, ANID BASAL project FB210003, FONDECYT Regular grant 1210425, CASSACA grant CCJRF2105, and ANID+REC Convocatoria Nacional subvencion a la instalacion en la Academia convocatoria 2020 PAI77200076. MT acknowledges support from the Jenny and Antti Wihuri Foundation. This work has also made use of the University of Hertfordshire’s high-performance computing facility.
Appendix A Complete Radial Velocity and Stellar Activity Index Data
The following tables list all of the RVs and stellar activity index data utilized in this paper.
Table 8: Complete AAT data
JD RV RV_error
2451034.17836 -16.150 2.740
2451413.09591 -22.900 2.940
2451683.28594 -10.290 2.810
2451743.20344 30.660 3.250
2451767.12176 18.660 2.330
2451828.93670 11.580 4.030
2452062.28449 -9.140 2.350
2452092.18604 4.530 2.540
2452130.09003 20.530 2.830
2452151.06087 16.690 2.690
2452151.99385 12.150 2.330
2452152.97460 16.090 2.160
2452154.05537 3.020 3.310
2452186.93709 -16.560 2.120
2452188.01333 -26.390 2.180
2452189.07521 -17.590 1.980
2452189.95958 -15.190 2.420
2452388.30750 4.590 2.450
2452477.16326 1.440 3.070
2452598.94574 -0.560 2.690
2452784.26723 24.630 1.840
2452859.14214 2.890 2.470
2452943.06869 5.370 2.540
2452946.99543 -3.200 2.790
2453214.19050 -14.330 2.250
2453215.18430 -10.380 1.900
2453216.22254 -14.000 1.730
2453243.24872 -20.500 2.940
2453244.14557 -9.770 2.240
2453245.14036 -22.760 2.640
2453245.93542 -9.930 2.700
2453282.01051 3.900 5.840
2453489.27774 -12.640 1.950
2453573.18022 37.850 2.800
2453664.98472 13.860 1.950
2453945.22713 19.400 1.360
2454012.98659 -22.590 1.490
2454038.99785 -8.610 3.210
2454254.21331 -22.420 2.130
2454338.22561 21.420 2.370
2454371.09020 -1.740 1.700
2455103.01798 -0.130 2.050
2455399.21234 12.200 2.110
2455430.18814 -0.880 2.560
2455845.97594 12.180 2.830
2456467.27664 -14.020 2.400
2456935.06220 11.780 2.990
2456969.95622 -18.000 3.300
Table 9: Complete TERRA data
JD RV RV_error S-index FWHM BIS
2453265.69209 -22.310 0.817 0.164 8.465 22.481
2453269.66244 -17.369 1.281 0.163 8.468 19.289
2453861.92676 -34.401 0.687 0.162 8.461 20.829
2453862.92224 -34.217 0.643 0.162 8.459 17.286
2453863.94041 -34.169 0.634 0.163 8.466 19.206
2453864.88194 -34.095 0.691 0.164 8.465 19.668
2453865.92725 -30.038 0.553 0.164 8.457 20.273
2453866.90195 -31.316 0.777 0.164 8.464 19.119
2453867.92542 -35.551 0.762 0.162 8.464 22.462
2453868.91623 -34.276 0.702 0.162 8.463 18.725
2453869.88794 -36.017 0.893 0.161 8.463 18.353
2453870.87586 -35.859 0.427 0.159 8.463 17.593
2453882.92017 -35.249 0.824 0.160 8.465 20.386
2453883.90239 -33.331 0.622 0.160 8.460 17.748
2453886.91959 -34.368 0.611 0.161 8.466 21.496
2453887.90246 -28.890 0.698 0.161 8.467 16.819
2454254.91746 -37.187 1.115 0.161 8.469 18.714
2454255.82308 -37.086 0.891 0.163 8.463 19.590
2454257.90216 -30.784 1.004 0.165 8.462 22.762
2454258.93746 -31.501 1.006 0.167 8.472 18.799
Table 10: Complete PFS data
JD RV RV_error S-index
2455847.60360 -7.090 2.048 0.159
2456094.81496 -15.712 1.943 0.147
2456141.70123 7.073 1.676 0.142
2456552.63221 18.083 2.075 0.163
2456867.78785 -20.282 1.960 0.147
2456876.82515 -23.179 1.897 0.144
2457199.84220 13.107 1.641 0.142
2457206.84952 3.411 1.813 0.140
2457259.72419 -7.563 2.018 0.153
2457321.58081 23.976 1.769 0.144
2457327.62340 22.557 1.601 0.144
2457529.89692 -10.381 1.543 0.143
2457554.89578 -1.735 1.763 0.146
2457616.64569 -29.852 1.633 0.140
2457620.64176 -24.033 2.066 0.339
2458266.84871 -14.930 1.449 0.150
2458272.91464 -11.629 1.389 0.145
2458293.80551 -7.507 1.776 0.148
2458297.82056 -2.482 1.591 0.145
2458298.91657 0.899 1.614 0.144
2458355.65345 26.318 1.910 0.157
2458355.65588 21.484 1.864 0.158
2458355.65830 27.853 1.945 0.157
2458625.87440 4.666 1.677 0.163
2458625.87721 6.568 1.821 0.162
2458625.88005 3.032 1.769 0.162
2458674.82066 -16.397 1.823 0.152
2458674.82665 -13.872 1.755 0.152
2458685.76083 -7.206 1.650 0.152
2458685.76531 -7.692 1.759 0.157
2458738.60957 16.696 1.888 0.159
2458738.61344 17.112 1.899 0.158
2459152.57287 -3.915 1.665 0.150
2459152.57582 -5.512 1.626 0.153
2459152.57882 -0.662 1.668 0.153
2459359.89486 3.947 1.529 0.158
2459359.89679 0.659 1.597 0.158
2459359.89874 4.639 1.687 0.156
2459359.90067 3.909 1.690 0.157
2459449.74696 -16.079 1.871 0.160
2459449.75193 -15.890 1.990 0.163
2459475.63281 -10.275 1.766 0.157
2459475.63674 -13.527 1.810 0.157
2459508.64032 1.264 1.723 0.161
2459508.64415 2.893 1.742 0.160
Table 11: Complete TERRA2 data
JD RV RV_error S-index FWHM BIS
2457559.81282 9.378 1.172 0.157 8.496 30.035
2457559.81885 8.200 1.215 0.157 8.497 27.733
2457559.82494 8.462 1.326 0.158 8.497 23.558
2457560.70802 8.305 0.900 0.156 8.494 25.718
2457560.71411 7.898 1.010 0.157 8.498 25.647
2457560.72037 10.108 1.042 0.156 8.496 25.686
2457561.77437 11.317 0.733 0.157 8.499 27.346
2457561.78040 10.581 0.969 0.156 8.501 30.131
2457561.78661 10.048 1.067 0.156 8.497 26.460
2457609.72818 -17.828 1.046 0.156 8.498 23.096
2457609.73450 -18.772 0.910 0.157 8.500 22.999
2457609.74047 -15.236 0.855 0.156 8.504 23.118
2457611.72416 -17.658 0.954 0.156 8.495 21.300
2457611.73013 -18.175 0.902 0.157 8.497 27.250
2457611.73611 -17.363 0.892 0.156 8.498 24.630
2457611.74231 -17.838 0.982 0.157 8.491 23.433
2458004.65674 -21.669 0.756 0.157 8.498 27.683
2458004.66278 -23.562 0.936 0.157 8.494 27.091
2458004.66910 -21.507 0.976 0.156 8.501 25.370
2458084.53776 12.779 0.997 0.157 8.494 24.403
2458084.54090 11.632 1.171 0.157 8.503 25.691
2458084.54403 13.232 1.040 0.159 8.494 26.135
2458084.54721 14.627 0.872 0.157 8.504 28.006
2458084.55028 12.451 1.178 0.159 8.498 27.758
2458084.55346 14.958 1.261 0.159 8.498 27.041
2458085.53276 10.846 1.359 0.157 8.498 27.531
2458085.53537 14.197 1.562 0.154 8.494 19.238
2458085.53804 13.877 1.916 0.152 8.501 28.371
2458085.54086 9.780 1.500 0.156 8.493 29.324
2458085.54346 13.102 1.461 0.156 8.493 25.662
2458224.89687 26.569 0.590 0.158 8.500 22.623
2458224.90303 26.603 0.939 0.158 8.499 22.645
2458224.90959 25.062 0.951 0.158 8.499 24.564
2458225.86378 26.246 0.958 0.158 8.495 24.616
2458225.86989 28.860 1.237 0.158 8.486 25.949
2458225.87611 22.589 1.429 0.158 8.499 35.029
2458241.87799 9.264 0.839 0.158 8.489 28.001
2458241.88414 8.849 0.907 0.160 8.492 26.270
2458241.89029 8.453 0.905 0.160 8.496 26.826
2458249.85897 -6.845 1.002 0.157 8.492 26.792
2458249.86391 -5.844 0.809 0.156 8.492 26.581
2458249.86890 -3.477 0.856 0.159 8.491 23.333
2458251.91615 -6.078 1.258 0.158 8.489 26.863
2458251.92259 -8.067 1.239 0.159 8.488 31.964
2458251.92868 -4.637 1.020 0.157 8.491 26.071
2458252.92118 -5.254 1.144 0.158 8.496 28.363
2458252.92732 -6.066 1.129 0.158 8.494 28.735
2458252.93318 -5.229 0.807 0.159 8.492 26.200
2458318.74664 5.564 2.200 0.158 8.497 25.277
2458318.75416 6.612 1.571 0.160 8.498 21.977
2458318.76112 8.342 2.156 0.159 8.507 25.269
2458324.79862 10.921 1.133 0.161 8.499 28.935
2458324.80482 11.625 0.962 0.160 8.496 22.548
2458324.81096 10.200 1.140 0.158 8.497 25.251
2458325.84981 10.026 1.118 0.159 8.501 27.630
2458325.85590 10.939 1.066 0.157 8.506 29.735
2458325.86209 11.762 0.962 0.158 8.501 25.440
2458325.86806 10.217 0.990 0.157 8.498 26.114
2458325.87420 10.407 1.210 0.158 8.506 26.382
2458384.63240 -10.514 1.172 0.155 8.497 27.505
2458384.63599 -10.399 0.944 0.155 8.497 22.623
2458384.63955 -13.060 1.181 0.158 8.495 28.033
2458385.57186 -10.301 1.228 0.157 8.492 30.020
2458385.57515 -11.608 1.316 0.157 8.493 16.960
2458385.57842 -13.690 1.403 0.158 8.495 27.307
2458386.77532 -9.965 0.988 0.157 8.489 24.726
2458386.77857 -8.429 1.083 0.157 8.496 26.359
2458386.78183 -11.937 1.226 0.156 8.495 17.562
2458407.63040 -15.514 0.888 0.156 8.495 29.768
2458407.63675 -14.502 1.343 0.156 8.497 23.465
2458407.64275 -12.674 1.087 0.156 8.494 25.263
2458438.49764 1.811 0.976 0.159 8.493 27.538
2458438.50378 -0.626 0.935 0.159 8.496 27.039
2458438.50998 -0.327 1.002 0.159 8.497 24.048
2458440.50190 1.509 1.057 0.158 8.535 26.228
2458440.50793 2.065 0.613 0.157 8.498 26.406
2458440.51408 0.000 0.846 0.158 8.498 24.668
2458452.52558 5.358 0.836 0.157 8.506 27.655
2458452.52943 4.142 0.963 0.160 8.495 22.377
2458452.53331 3.154 1.239 0.159 8.501 27.534
2458455.50983 4.563 1.374 0.158 8.500 22.960
2458455.51372 4.966 1.258 0.161 8.498 22.328
2458455.51782 7.583 1.679 0.160 8.497 22.885
2458590.93608 4.165 0.483 0.161 8.482 27.989
2458599.92354 13.662 0.965 0.164 8.496 27.231
2458599.92976 14.661 0.793 0.161 8.500 29.973
2458635.90159 -5.868 1.619 0.165 8.501 32.074
2458635.90788 -4.601 1.369 0.162 8.494 24.752
2458635.91398 -3.983 1.480 0.163 8.499 29.327
2458666.81934 -10.873 1.339 0.167 8.517 25.558
2458666.82548 -9.076 1.271 0.163 8.519 17.495
2458666.83146 -10.298 1.340 0.166 8.522 27.470
2458677.85869 -6.515 1.115 0.164 8.493 29.236
2458677.86491 -8.033 1.084 0.162 8.496 26.649
2458677.87089 -9.180 1.212 0.163 8.495 28.461
2458695.67418 -0.342 0.976 0.165 8.510 24.550
2458726.60270 13.250 1.448 0.167 8.498 32.783
2458726.60885 12.321 1.424 0.165 8.504 27.081
2458726.61505 7.810 1.675 0.167 8.495 30.461
2458734.62837 19.746 1.065 0.164 8.514 21.242
2458734.63424 20.831 1.207 0.166 8.506 33.235
2458734.64056 19.867 1.111 0.164 8.516 25.962
2458754.67141 11.183 1.096 0.165 8.515 23.612
2458754.67743 14.470 1.089 0.165 8.523 23.461
2458754.68351 13.521 0.913 0.165 8.516 28.589
2458755.48311 10.043 0.907 0.164 8.522 22.886
Appendix B Split RV GLS Periodograms
Figure 10 shows the GLS Periodogram of the Low S-index dataset, while Figure 11 shows the GLS Periodogram of the High S-index dataset.
Figure 10: Top Panel: Original GLS periodogram of the combined and mean-subtracted Low S-index RV data. Second Panel: GLS periodogram after fitting and removing the first signal. Third Panel: GLS periodogram after fitting and removing both the first and second signals. Bottom Panel: GLS periodogram after manually fitting the third signal and removing all three previously detected signals. The blue, dotted horizontal lines represents the 0.1% False Alarm Probability (FAP) level estimated from 5000 bootsrap resamplings.
Figure 11: Top Panel: Original GLS periodogram of the combined and mean-subtracted High S-index RV data. Second Panel: GLS periodogram after fitting and removing the first signal. The blue, dotted horizontal lines represents the 0.1% False Alarm Probability (FAP) level estimated from 5000 bootsrap resamplings.